At time 13, observed was (1,3,7)

At time 84, observed was (23, -5, 10)

At time 50 , What are the positions observed?
Answer #1

Change in position from 13 to 84 is (22,-8,3) so velocity will be cange in postion divided by change in time and here time t=71 sec

So V = (22/71, -8/71,3/71)

Now at 50 change in time is 50-13= 37

So position at 50 = (1,3,7) +V×37

= (1+11.46, 3-4.17,7+1.56)

= (12.46,-1.17,8.56)

Solved